A garden measuring 12 meters by 16 meters is to have a pedestrian pathway installed around it, increasing the total area to 285 square meters. What will be the width of the pathway?
Answer by scott8148(6628)

(2w + 12)(2w + 16) = 285

4w^2 + 56w + 192 = 285 ___ 4w^2 + 56w - 93 = 0

factoring ___ (2w + 31)(2w - 3) = 0

2w + 31 = 0 ___ 2w = -31 ___ w = -31/2 ___ negative value not realistic

2w - 3 = 0 ___ 2w = 3 ___ w = 3/2